# On the Interchangeability of Positional Embeddings in Multilingual Neural Machine Translation Models

### v1

@@ Abstract
Standard Neural Machine Translation (NMT) models have traditionally been trained with Sinusoidal Positional Embeddings (PEs), which are inadequate for capturing long-range dependencies and are inefficient for long-context or document-level translation. In contrast, state-of-the-art large language models (LLMs) employ relative PEs, demonstrating superior length generalization. This work explores the potential for efficiently switching the Positional Embeddings of pre-trained NMT models from absolute sinusoidal PEs to relative approaches such as RoPE and ALiBi. Our findings reveal that sinusoidal PEs can be effectively replaced with RoPE and ALiBi with negligible or no performance loss, achieved by fine-tuning on a small fraction of high-quality data. Additionally, models trained without Positional Embeddings (NoPE) are not a viable solution for Encoder-Decoder architectures, as they consistently under-perform compared to models utilizing any form of Positional Embedding. Furthermore, even a model trained from scratch with these relative PEs slightly under-performs a fine-tuned model, underscoring the efficiency and validity of our hypothesis.
On the Interchangeability of Positional Embeddings in
Multilingual Neural Machine Translation Models
Varun Gumma  Pranjal A. Chitale  Kalika Bali Microsoft Corporation
@@ 1 Introduction
Neural Machine Translation (NMT) models have become integral to a broad spectrum of applications across various Natural Language Processing (NLP) tasks, particularly in generating synthetic data for training multilingual Large Language Models (LLMs). Despite their widespread use, state-of-the-art NMT systems often encounter sequence length constraints due to their reliance on Sinusoidal Positional Embeddings. Recent advancements have introduced alternative methods for embedding positional information, such as Rotary Positional Embeddings (RoPE) Su et al. (2024) and Attention with Linear Biases (ALiBi) Press et al. (2022), which demonstrate improved extrapolation capabilities for longer contexts. However, retraining existing models with these new embeddings is often computationally infeasible. This raises a critical question: Can we alter the Positional Embeddings in pre-trained models while preserving their performance? To the best of our knowledge, no prior study has examined the feasibility of swapping out Positional Embeddings in pre-trained transformer models.
This study addresses this question by investigating the feasibility of substituting Positional Embeddings in pre-trained NMT models. Additionally, inspired by prior works (Haviv et al., 2022; Kazemnejad et al., 2023), we explore whether Positional Embeddings are necessary in Encoder-Decoder models and whether these models can recover performance through fine-tuning after the removal of explicit positional information.
Figure 1: An basic overview of switching out Positional Embeddings
We conduct a thorough investigation into both parameter-efficient fine-tuning (PEFT) methods and full fine-tuning (FFT) strategies to restore performance and adapt pre-trained models to new Positional Embeddings. Our results demonstrate that it is indeed possible to substitute Positional Embeddings post-hoc without compromising model performance. Additionally, we find that adding extra parameters specifically to the self-attention block is sufficient for effective adaptation. This approach not only preserves sentence-level performance but also improves extrapolation capabilities for longer contexts when alternative Positional Embeddings, such as RoPE or ALiBi, are introduced.
In the remainder of this paper, we make the following contributions:
  • We provide a comprehensive analysis of the feasibility of modifying the Positional Embeddings of pre-trained transformer models post-hoc while maintaining comparable performance levels.
  • We demonstrate that these models can be adapted in a parameter-efficient manner by simply incorporating additional parameters into the self-attention block.
  • Our findings show that RoPE and ALiBi exhibit equivalent performance, while NoPE significantly degrades performance, highlighting it as a sub-optimal design choice for encoder-decoder transformer models.
  • We establish that post-hoc modification of Positional Embeddings achieves performance comparable to that of a model trained from scratch with the respective Positional Embeddings.
@@
  • Lastly, we open-source111 https://github.com/VarunGumma/fairseq our framework to facilitate the efficient training and fine-tuning of long-context machine translation models.
@@ 2 Related Works
Transformer-based models require an explicit Positional Embedding information as all the timesteps are processed parallely, and hence ther is a need for explicit positional. To this end, Vaswani et al. (2017) propose an empirical absolute Positional Embedding approach following a fourier series with alternating sine and cosine terms Kernes (2024). Devlin et al. (2019) introduce trainable Positional Embeddings which are learnt along with the other parameters of the model, however they have not shown a significant performance difference from prior static versions. Press et al. (2022) discuss ALiBi, a simple relative Positional Embedding strategy by adding the relative distance between tokens to the attention matrix and show that it can help the model generalize to longer sequence during inference even though trained on shorter sequences. RoPE Su et al. (2024) introduces the concept of rotating vectors according to their positions to explicitly encode relative position dependency in the self-attention formulation. Recent works Sun et al. (2023); Peng et al. (2023); Ding et al. (2024); Ostmeier et al. (2024) have also provided fixes to adapt RoPE to much longer sequences. Kazemnejad et al. (2023) find that no Positional Embedding can represents both relative and absolute strategies, and that it elicits high length generalizability. We refer the readers to Wang and Chen (2020) for a visual understanding of Positional Embeddings methods.
@@
Houlsby et al. (2019); Bapna and Firat (2019); Hu et al. (2022) deals with the concept of updating only a certain subset of parameters of the network, to efficiently generalize the model to new tasks or domains, instead of tuning all parameters, which might lead to catastrophic forgetting or over-fitting in large models. Low-Rank Adaptation (LoRA) Hu et al. (2022) is currently the most popular parameter efficient finetuning strategy, which hypothesizes that the update to weights are low-rank, and injects small trainable parameters directly into the pre-trained weights to mimic the low-rank decomposed updates. An extension, Weight-Decomposed Low-Rank Adaptation (DoRA) Liu et al. (2024) decomposes the pre-trained weight into a magnitude and direction, and employs LoRA only for directional updates to efficiently minimize the number of trainable parameters.
@@ 3.1 Model
Our experimentation focuses on standard Encoder-Decoder transformer models. While several massively multilingual models, such as NLLB (Costa-jussà et al., 2022) and its distilled variants, have been released, our experimentation requires the availability of long-context or document-level training and evaluation datasets. Consequently, we focus on language-family specific models that allow for more controlled experimentation. For our experiments, we chose the distilled variants of IndicTrans2 Gala et al. (2023). A key difference is that the distilled models released in Gala et al. (2023) we used were not trained with BackTranslation Sennrich et al. (2016) data (BPCC-BT), which we incorporate in our experimental setup. We experimented with models based on RoPE, ALiBi, and NoPE by replacing the sinusoidal Positional Embedding in the pre-trained model with the corresponding relative Positional Embedding. This replacement led to a significant performance drop, prompting us to fine-tune the models using high-quality data Mohiuddin et al. (2022); Gumma et al. (2023) to study their recovery in each setting.
@@ 3.2.1 Train set
Fine-tuning
For maximum performance gains for our models, we choose a high-quality mixture of seen and unseen distributions for our training data. To this end, for the seen distribution, we select the BPCC-Seed, which is a high-quality human-verified subset BPCC. This subset consists of X pairs, and has shown to be effective in improving model performance Gala et al. (2023). As for the unseen distribution, we sample the top 150K pairs for each language from the BackTranslation (BPCC-BT) subset of BPCC based on the provided LASER Costa-jussà et al. (2022) or LaBSE Feng et al. (2022) cosine-similarity scores. In case similarity scores are unavailable for a translation pair, a random sample was chosen. Additionally, for the unseen data, we also include the Asian Language Treebank [^2] https://www2.nict.go.jp/astrec-att/member/mutiyama/ALT (ALT) Corpus Riza et al. (2016) which contains long-context translation pairs for Hindi and Bengali, and the CoPara Aligned corpus E et al. (2023), which contains document-level translation pairs for 4 Dravidian languages, Telugu, Kannada, Malayalam and Tamil. Overall, our train set consists of 4.45M and 3.46M translation pairs across 22 languages in the X-En and En-X directions respectively.
Training
We combine the fine-tuning data prepared above with the training data of the IT2 distilled models as described in Gala et al. (2023), to create 114.09M and 113.10M translation pairs for the X-En and En-X directions, respectively.
@@ 3.2.2 Dev/Test sets
Most existing benchmarks for Indian languages are composed of sentence-level test sets, which may not adequately reflect long-context performance or extrapolation capabilities. To address this, we construct document-level development and test sets using Flores [^3] https://huggingface.co/datasets/facebook/Flores (Costa-jussà et al., 2022) and IN22-Conv [^4] https://huggingface.co/datasets/ai4bharat/IN22-Conv (Gala et al., 2023). In this setup, sentences are grouped within the same context window, making them suitable for document-level evaluation. For Flores, we group sentences by metadata, including URL, domain, and topic headers, to map sentence-level pairs into document-level pairs, where each document-level unit consists of three sentences. This method is viable since this benchmark was created by professional translators who translate three consecutive sentences from every selected context window, as described by Goyal et al. (2022). In case of IN22-Conv, the dataset was originally created by translating entire conversations from English into the target languages by professional translators, and used for sentence-level evaluation by Gala et al. (2023). We use metadata provided to trace back the conversations and treat the entire conversation as a sequence for long-context evaluation. As a result, we obtain 289 merged sentence pairs per language, with a total of 6,380 pairs.
3.3 Metrics and Evaluation
For consistency, we use the ChrF++ score (Popović, 2017) as the primary metric, and follow the same evaluation practices as outlined in Gala et al. (2023) by normalizing the indic texts using indic_nlp_library [^5] https://github.com/anoopkunchukuttan/indic_nlp_library. For brevity, we report the aggregate score across all the languages in the respective test set, which are denoted in Table 1.
@@ 3.3 Metrics and Evaluation
Table 1: List of languages in each test set.
@@ 3.4 Finetuning Strategies
In our fine-tuning experiments, we swap the sinusoidal Positional Embedding module with a RoPE, ALiBi or NoPE module, and fine-tune the model as efficiently as possible from that point to analyse the performance recovery.
@@
First, we choose a full-fine-tuning (FFT) approach which tweaks all the 211.78M parameters of the model. This would allow the embeddings, layer-normalizations, and the rest of the linear layers666self-attention, cross-attention, feed-forward modules to be tuned as per the new relative Positional Embedding.
@@ 4 Experiments
We use the fairseq [^7] Our variant implements, LoRA, RoPE, ALiBi, and a slightly efficient SDPA as per the requirements of this work framework Ott et al. (2019) for all of our experiments. All the finetuning experiments were conducted on 4 $\times$ A100 80Gb GPUs, the train-from-scratch experiments on 8 $\times$ H100 80Gb GPUs, and every model was trained till convergence with early-stopping. The exact hyperparameter sets are mentioned in Tables 2 and 3.
@@
Checkpoint 1000 steps, BLEU @ $k=5$
@@
Table 2: Default hyperparameter settings for finetuning. The cells with two numbers present the values for FFT/LoRA. LoRA finetuning acts on very specific modules, and adds minimal trainable parameters, while keeping the rest frozen. This acts like an implicit regularization, and hence we do not add any explicit strategies like label-smoothing and dropout to the model in this setup.
Hyperparameter Value
Global Max Tokens 256K
Optimizer Adam
LR scheduler inverse_sqrt
Gradient clip norm 1.0
Adam betas (0.9, 0.98)
Checkpoint 1000 steps, BLEU @ $k=5$
Patience 10
Max Positions (4096, 4096)
Warmup steps 4000
Learning rate 7e-4
Dropout 0.2
Label smoothing 0.1
FP16 True
Table 3: Default hyperparameter settings for training from scratch
4.1 Baselines
We compare our hypothesis against several baselines as outlined below. First, we examine a direct swapping approach, where models with sinusoidal Positional Embeddings are altered to not use Positional Embeddings (NoPE) or subject to alternate Positional Embeddings like RoPE, ALiBi, and their performance is evaluated. Additionally, we consider fine-tuning these models under three conditions outlined in section 3.4. Finally, to estimate a bound for each Positional Embedding method without swapping, we train transformer models from scratch using Sinusoidal, RoPE, ALiBi, and NoPE.

### v2

# Towards Inducing Document-Level Abilities in Standard Multilingual Neural Machine Translation Models

@@ Abstract
Neural Machine Translation (NMT) models have traditionally used Sinusoidal Positional Embeddings (PEs), which often struggle to capture long-range dependencies and are less efficient for handling extended context or document-level translation tasks. This work addresses the challenge of transitioning pre-trained NMT models from absolute sinusoidal PEs to relative PEs, such as Rotary Positional Embeddings (ROPE) and Attention with Linear Biases (ALIBI), without compromising performance. We demonstrate that parameter-efficient fine-tuning, using only a small amount of high-quality data, can successfully facilitate this transition. Experimental results indicate that switching from sinusoidal to relative PEs results in competitive translation quality on sentence-level evaluation benchmarks. Additionally, models trained with ROPE consistently outperform those using ALIBI and Sinusoidal PEs on document-level benchmarks across both string-based metrics and qualitative evaluations. Moreover, we find that a small amount of long-context data in a few languages is sufficient for cross-lingual length generalization, thereby inducing long-context capabilities.
Towards Inducing Document-Level Abilities in Standard Multilingual Neural Machine Translation Models
Varun Gummathanks:  Equal contribution  Pranjal A. Chitale11footnotemark: 1  Kalika Bali Microsoft Corporation
@@ 1 Introduction
Neural Machine Translation (NMT) models have become critical tools in many Natural Language Processing (NLP) applications, especially for document-level translation. While NMT models perform well on sentence-level tasks, they often struggle with longer sequences, where preserving contextual coherence across sentences is essential. This limitation is largely due to their reliance on Absolute Positional Embedding (PE) methods, such as Sinusoidal PEs, which can limit the model’s ability to extrapolate effectively to longer contexts. Recent advancements in Relative PEs, such as Rotary Positional Embeddings (ROPE) and Attention with Linear Biases (ALIBI), have demonstrated superior performance in handling longer sequences. These relative positional embedding techniques enable models to generalize better to long contexts, making them well-suited for document-level translation tasks. However, most state-of-the-art NMT models were originally trained using sinusoidal embeddings, and retraining them with newer methods like ROPE or ALIBI is computationally expensive.
This presents a crucial question: Can we modify the positional embeddings in pre-trained NMT models to improve long-context translation without significant retraining or performance degradation? Addressing this problem would enable NMT models to translate long documents more effectively, improving translation coherence and quality across paragraphs and sections. Moreover, long-context NMT models could have broader applications, such as generating synthetic data for multilingual pre-training. Such data is often needed in scenarios where natural multilingual corpora are limited, especially for low-resource languages. However, current sentence-level NMT systems are suboptimal for this task as they lack the capacity to capture broader contextual dependencies in long documents.
In this work, we investigate whether it is possible to replace the positional embeddings in pre-trained NMT models with relative positional embeddings post-hoc. Our study explores both parameter-efficient fine-tuning (PEFT) methods and full fine-tuning (FFT) strategies to restore model performance while enabling long-context generalization.
In the remainder of this paper, we present the following contributions:
  • We conduct a comprehensive study on modifying positional embeddings in pre-trained transformer models post-hoc, aiming to retain performance on existing tasks while enabling long-context capabilities.
  • Experiments show that ROPE and ALIBI perform similarly on sentence-level benchmarks, but ROPE outperforms both ALIBI and SINE on document-level tasks, based on string-based evaluation metrics, which was then further validated by qualitative evaluation.
  • Moreover, fine-tuning with ROPE using minimal long-context data across a few languages demonstrates effective cross-lingual length generalization, even without extensive long-context training data.
@@
  • Lastly, we open-source111 https://github.com/VarunGumma/fairseq our framework to support the efficient training and fine-tuning of long-context machine translation models.
Our findings indicate that positional embeddings in pre-trained NMT models can be modified and adapted without compromising performance, to efficiently induce long-context capabilities.
@@ 2 Related Works
Transformer-based models require an explicit Positional Embedding (PE) information as all the timesteps are processed parallely, and hence there is a need for explicit positional. To this end, Vaswani et al. (2017) propose an empirical absolute PE approach following a fourier series with alternating sine and cosine terms. Vaswani et al. (2017) also experiment with Learnable PEs, however they do not show a significant performance difference as compared to absolute Sinusoidal Embeddings. Later works Rosendahl et al. (2019); Shaw et al. (2018) have demonstrated the advantages of relative PEs in transformer. Press et al. (2022) discuss ALIBI, a simple relative PE strategy achieved by adding the relative distance between tokens to the attention matrix and show that it can help the model generalize to longer sequence during inference even though trained on shorter sequences. Rotary Positional Embeddings (ROPE) Su et al. (2024) introduces the concept of rotating vectors according to their positions to explicitly encode relative position dependency in the self-attention formulation. Recent works Sun et al. (2023); Peng et al. (2023); Ding et al. (2024); Ostmeier et al. (2024) have also provided fixes to adapt ROPE to much longer sequences. In this work, we restrict ourselves to the standard variant of ROPE however, our method is generalizable to any other variant as well. In this study, we focus on the standard variant of RoPE, but our approach would be generalizable to other variants as well.
@@
Houlsby et al. (2019); Bapna and Firat (2019); Hu et al. (2022) deals with the concept of updating only a certain subset of parameters of the network, to efficiently generalize the model to new tasks or domains, instead of tuning all parameters, which might lead to catastrophic forgetting or over-fitting in large models. Low-Rank Adaptation (LORA) Hu et al. (2022) is currently the most popular parameter efficient finetuning strategy, which hypothesizes that the update to weights are low-rank, and injects small trainable parameters directly into the pre-trained weights to mimic the low-rank decomposed updates.
Evaluation metrics for Machine Translation
Evaluation in machine translation is a well-explored area, with string-based metrics like BLEU (Papineni et al., 2002), ChrF (Popović, 2015), and ChrF++ (Popović, 2017) being the most commonly used. However, prior studies (Kocmi et al., 2021) have pointed out the limitations of BLEU, noting its bias towards exact matches, which makes it less suitable for languages with flexible word order. Trained metrics, such as COMET (Rei et al., 2020, 2022) and BLEURT (Sellam et al., 2020), utilize a regressor on top of an encoder to predict translation quality. These metrics have been shown to correlate more closely with human judgments, outperforming traditional string-based measures (Kocmi et al., 2021). However, their performance can vary significantly across different languages, and may not always be reliable for Indian languages (Gala et al., 2023).
LLM-based Evaluators for Machine Translation
Lately, LLMs are emerging as general-purpose evaluators in Multilingual scenarios Hada et al. (2024b, a), exhibiting superior correlations with human judgment. Similar approaches have also been applied to the MT task as well. GEMBA (Kocmi and Federmann, 2023b) was one of the initial attempts to use LLMs to evaluate machine translation outputs. Follow up works like EAPROMPT (Lu et al., 2024), AutoMQM (Fernandes et al., 2023), GEMBA-MQM (Kocmi and Federmann, 2023a) have used LLMs to evaluate translations in a Multidimensional Quality Metrics (MQM) framework (Lommel et al., 2014; Freitag et al., 2021), by prompting them to identiy and penalize erroneous spans of text offering more interpretable and standardized evaluation. These appraoches demonstrate superior correlation with human judgements when compared to other model-based metrics like COMET, BLEURT. Given that LLMs typically support longer context lengths, they are well-suited for evaluating tasks involving extended input sequences. Consequently, inspired by GEMBA-MQM (Kocmi and Federmann, 2023a) we employ GPT-4o as an evaluator to investigate the qualitative differences between these baseline models.
@@ 3.1 Model
Our experimentation focuses on standard Encoder-Decoder transformer models. While several massively multilingual models, such as NLLB (Costa-jussà et al., 2022) and its distilled variants, have been released, our experimentation requires the availability of long-context or document-level training and evaluation datasets. Consequently, we focus on language-family specific models that allow for more controlled experimentation. For our experiments, we chose the distilled variants of IndicTrans2 Gala et al. (2023); Gumma et al. (2023). A key difference is that the distilled models released in Gala et al. (2023) we used were not trained with BackTranslation Sennrich et al. (2016) data, which we incorporate in our experimental setup. We experimented with models based on ROPE amd ALIBI by replacing the sinusoidal Positional Embedding in the pre-trained model with the corresponding relative Positional Embedding. This replacement led to a significant performance drop, prompting us to fine-tune the models using high-quality data Mohiuddin et al. (2022); Gumma et al. (2023); Gala et al. (2023) to study their recovery in each setting.
@@ 3.2.1 Train set
For maximum performance gains for our models, we choose a high-quality mixture of seen and unseen distributions for our training data. To this end, for the seen distribution, we select the BPCC-Seed, which is a high-quality human-verified subset of BPCC (Bharat Parallel Corpus Collection), and has shown to be effective in improving model performance Gala et al. (2023). As for the unseen distribution, we sample the top 150K pairs for each language from the BackTranslation (BPCC-BT) subset of BPCC based on the provided LASER Costa-jussà et al. (2022) or LaBSE Feng et al. (2022) cosine-similarity scores222whichever was available for that language. In case similarity scores are unavailable for a translation pair, a random sample was chosen. Additionally, for the unseen data, we also include the Asian Language Treebank [^3] https://www2.nict.go.jp/astrec-att/member/mutiyama/ALT (ALT) Corpus Riza et al. (2016), and the CoPara Aligned corpus E et al. (2023), which contains document-level translation pairs for 4 Dravidian languages, Telugu, Kannada, Malayalam and Tamil. The overall statistics of the dataset is available in table 1.
BPCC
Seed
BPCC
BT
ALT CoPara Total
En-Indic 1.41 2.76 0.04 0.02 4.23
Indic-En 1.41 3.75 0.04 0.02 5.22
Table 1: Number of training instances per subset (in Millions)
256-
512
512-
1024
1024-
2048
2048-
4096
Total
eng_Latn-
kan_Knda
325 66 4 4 399
eng_Latn-
mal_Mlym
313 75 4 3 395
eng_Latn-
tam_Taml
331 77 5 3 416
eng_Latn-
tel_Telu
321 71 5 4 401
Total 1290 289 18 14 1611
Table 2: Distribution of long-context data in the fine-tuning data mixture. Indic-En pairs also follow the same distribution.
@@ 3.2.2 Dev/Test sets
Most existing benchmarks for Indian languages are composed of sentence-level test sets, which may not adequately reflect long-context performance or extrapolation capabilities. To address this, we construct document-level development and test sets using Flores [^4] https://huggingface.co/datasets/facebook/Flores (Costa-jussà et al., 2022) and IN22-Conv [^5] https://huggingface.co/datasets/ai4bharat/IN22-Conv (Gala et al., 2023). In this setup, sentences are grouped within the same context window, making them suitable for document-level evaluation. For Flores, we group sentences by metadata, including URL, domain, and topic headers, to map sentence-level pairs into document-level pairs, where each document-level unit consists of three sentences. This method is viable since this benchmark was created by professional translators who translate three consecutive sentences from every selected context window, as described by Goyal et al. (2022). As a result, we obtain 286 merged sentence pairs per language, with a total of 5,720 pairs. In case of IN22-Conv, the dataset was originally created by translating entire conversations from English into the target languages by professional translators, and used for sentence-level evaluation by Gala et al. (2023). We use metadata provided to trace back the conversations and treat the entire conversation as a sequence for long-context evaluation. This dataset has 44 conversations per language and covers 22 languages. The token length distribution of the synthesized IN22-Conv test set is shown in Figure 1. IN22-Conv serves as an accurate representation of a true document-level evaluation set due to its contextual dependencies across multiple turns in the conversation. Additionally, it supports all the languages under consideration, making it the primary benchmark for document-level evaluation.
Figure 1: Distribution of source token lengths for the Indic side of conversations in the IN22-Conv Doc-Level test set tokenized using the IndicTrans2 tokenizer. The red line is the average across all languages.
@@
Table 3: List of languages in each test set. The top-12 languages are highlighted.
3.3 Metrics and Evaluation
For consistency, we use the ChrF++ score (Popović, 2017) as the primary string-based metric, and follow the same evaluation practices as outlined in Gala et al. (2023) by normalizing the indic texts using indic_nlp_library [^6] https://github.com/anoopkunchukuttan/indic_nlp_library. For brevity, we report the aggregate score across all the languages in the respective test set, which are denoted in Table 3. For GPT-4o-based evaluation, we use the prompt template from GEMBA-MQM (Kocmi and Federmann, 2023a) as a starting point and make slight adjustments to prompt the model to additionally provide an implicit quality score between 0 and 1, in addition to identifying the error spans follwing MQM guidelines. Another notable difference is that, since we are evaluating long-context translations, including few-shot examples would be inefficient due to the limited context length. Therefore, we opt for a zero-shot evaluation setup. We report both the implicit score as rated by GPT-4o as well as calculated MQM-score (based on the scoring guidelines outlined in Kocmi and Federmann (2023a). We start with an ideal score of 25, deducting points based on identified errors and their severity: critical errors are penalized by a factor of 5, major by 3, and minor by 1. The final score is normalized by dividing by 25, yielding a value between 0 and 1.
@@ 3.4 Finetuning Strategies
In our fine-tuning experiments, we swap the sinusoidal Positional Embedding module with a RoPE, or ALIBI module, and fine-tune the model as efficiently as possible from that point to analyse the performance recovery.
@@
First, we choose a full-fine-tuning (FFT) approach which tweaks all the 211.78M parameters of the model. This would allow the embeddings, layer-normalizations, and the rest of the linear layers777self-attention, cross-attention, feed-forward modules to be tuned as per the new relative Positional Embedding.
@@ 4 Experiments
We use the fairseq [^8] Our variant implements, LORA, ROPE, ALIBI, and a slightly efficient SDPA as per the requirements of this work framework Ott et al. (2019) for all of our experiments. All the finetuning experiments were conducted on 8 $\times$ H100 80Gb GPUs, and every model was trained till convergence with early-stopping. The exact hyperparameter sets are mentioned in Table 4.
@@
Checkpoint 1000 steps
Checkpoint Criteria BLEU @ $beam=5$
@@
Table 4: Default hyperparameter settings for finetuning. The cells with two numbers present the values for FFT/LoRA.